For each function,
i) identify the values of the parameters a, h, and k
ii) identify the transformations
iii) use transformations to graph the function and check that it is
correct with a table of values or a graphing calculator

f(x)=-x^2-2
1
Expert's answer
2020-12-01T09:11:09-0500

i). let f(a)= a2 + 5

a is the parameter

let f(h) = h2 - 8

h is the parameter

let f(k) = k2 + 6

k is a parameter

ii). for f(a) = a2 + 5

transformation is f(-a) = -a2 +5

for f(h) = h2 - 8

transformation is f(-h) = -h2 - 8

for f(k) = k2 + 6

transformation is f(-k) = -k2 + 6

iii). f(x) = -x2 - 2

If we transformation this function is

f(-x) = x2 - 2

The Graph is